Find the value of the determinants 4. a b a c d b) k 4- c) 5 2 7 8 d) 2-4- 8
Answer #1

4) Find the value q the determinant. a) 1 a b =) la bol = axd- bxc. = ad-bc. la bl=ad-bc. b) 1,5 -) = (x5)- (+1)x(1 = 10-(-)